Q1. Find the truth tables for the following statement forms:
1. p∨ ∼ q
2. p ∨ (q ∧ r )
3. (p ∨ q) ∧ (p ∨ r )
1
Expert's answer
2021-07-08T09:42:53-0400

Let us find the truth tables for the following statement forms:


1. "p\u2228 \u223c q"


"\ \begin{array}{||c|c||c|c||}
\hline \hline 
p & q & \sim q & p\lor \sim q \\
\hline\hline
0 & 0 & 1 & 1\\
\hline
0 & 1 & 0 & 0\\
\hline
1 & 0 & 1 & 1\\
\hline
1 & 1 & 0 & 1\\
\hline\hline
\end{array}"


2. "p \u2228 (q \u2227 r )"


"\ \begin{array}{||c|c|c||c|c||}
\hline \hline 
p & q & r & q\land r & p\lor ( q\land r) \\
\hline\hline
0 & 0 & 0 & 0 & 0\\
\hline
0 & 0 & 1 & 0 & 0\\
\hline
0 & 1 & 0 & 0 & 0\\
\hline
0 & 1 & 1 & 1 & 1\\
\hline
1 & 0 & 0 & 0 & 1\\
\hline
1 & 0 & 1 & 0 & 1\\
\hline
1 & 1 & 0 & 0 & 1\\
\hline
1 & 1 & 1 & 1 & 1\\
\hline\hline
\end{array}"


3. "(p \u2228 q) \u2227 (p \u2228 r )"


"\ \begin{array}{||c|c|c||c|c|c||}
\hline \hline 
p & q & r & p\lor q & p\lor r & (p\lor q)\land(p\lor r) \\
\hline\hline
0 & 0 & 0 & 0 & 0 & 0\\
\hline
0 & 0 & 1 & 0 & 1 & 0\\
\hline
0 & 1 & 0 & 1 & 0 & 0\\
\hline
0 & 1 & 1 & 1 & 1 & 1\\
\hline
1 & 0 & 0 & 1 & 1 & 1\\
\hline
1 & 0 & 1 & 1 & 1 & 1\\
\hline
1 & 1 & 0 & 1 & 1 & 1\\
\hline
1 & 1 & 1 & 1 & 1 & 1\\
\hline\hline
\end{array}"